caira12:

Cathy did a survey in four bird parks and wrote the following observations: Bird Park Survey Park Number of Green Birds Total Number of Birds P 18 82 Q 39 88 R 38 64 S 22 50 Based on the survey, which park had the greatest percentage of green birds? (5 points) Select one: a. Park P b. Park Q c. Park R d. Park S

Mercury:

for each park divide green birds/total # of birds and see which park gives you the greatest result

caira12:

so park Q?

Mercury:

check your calculations again, there's one that gives a bigger result than part Q

caira12:

park P

Mercury:

park P: 18/82 = 0.219 park Q: 39/88 = 0.443 park R: 38/64 = 0.594 park S: 22/50 = 0.44 which is the largest?

caira12:

ohhhhh Park S

Mercury:

notice how park R is 0.594 which is greater than 0.44 so park R is the greatest

caira12:

ohhhhhh